# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v4

@@ IX-B Radar Waveform KLD-based Optimisation
Figure 2: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using radar waveform optimisation: K-ROP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 2 illustrates the performance of both subsystems using the KLD-based radar waveform optimisation (K-ROP) technique. In Fig. 2.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ on the right axis and detection probability $P_{D}$ on the left axis versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance. By comparing the results here with the ZF-based design in Fig. 1, it can be noticed that there is a significant gain can be achieved using K-ROB. For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig. 1, demonstrating improved detection capabilities through K-ROP optimisation technique. For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system.
@@
Fig. 2.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ on the right axis and BER on the left axis versus $P_{\mathrm{R}}/N_{o}$. For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB suggests it has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced. The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted from Fig. 1. These results highlight the intricate relationship between radar detection capabilities and communication system performance.
@@ IX-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$. The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
@@
Figure 3: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using communication waveform optimisation: K-COP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 3 illustrates the performance of both subsystems using the KLD-based communication waveform optimisation (K-COP) technique. In Fig. 3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ (right axis) and detection probability $P_{D}$ (left axis) versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform. At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig. 1. K-COP demonstrates improved detection capabilities over non-optimisation but remains below the radar waveform optimisation performance achieved using K-ROP in Fig. 2, which is $\{98,89.2,99.9\}\%$.
@@
Fig. 3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ (right axis) and BER (left axis) versus $P_{\mathrm{R}}/N_{o}$. For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig. 1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs. K-COP optimisation technique provides significant performance gains over the non-optimised ZF scenario, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec. IX-B due to radar interference. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
@@ IX-D Integrated Waveform KLD-based Optimisation for ISAC
Figure 4: The ISAC system performance using the integrated waveform optimisation. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 4 illustrates the ISAC system performance under the KLD-based integrated waveform optimisation (K-IOP). In Fig. 4.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ (left axis) and $\mathrm{KLD}_{\mathrm{r}}$ (right axis) for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$. As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP. For instance, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, the achievable $\mathrm{KLD}_{\mathrm{r}}$ values for $\{\mathrm{T}_{1},\mathrm{T}_{2},\mathrm{T}_{3}\}$ are approximately $\{38.87,32.77,41.85\}$ bits, with corresponding detection probabilities $\{0.94,0.89,0.97\}$. Performance continues to rise at higher $\mathit{P}_{\mathrm{R}}/N_{o}$, achieving near-unity detection across all targets once $\mathit{P}_{\mathrm{R}}/N_{o}$ exceeds 22 dB.
@@
In Fig. 4.b, the communication subsystem performance is depicted via the BER (left axis) and $\mathrm{KLD}_{\mathrm{c}}$ (right axis) for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$. Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER. For example, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, $\mathrm{KLD}_{\mathrm{c}}$ values of $\{77.19,28.13,260.53\}$ bits are observed for $\{\mathrm{UE}_{1},\mathrm{UE}_{2},\mathrm{UE}_{3}\}$, with corresponding BERs of approximately $\{5.1\times 10^{-9},1.7\times 10^{-4},\lesssim 10^{-10}\}$. Beyond 14-16 dB, the BER for all UEs drops rapidly– often reaching effectively error-free communication– and $\mathrm{KLD}_{\mathrm{c}}$ grows significantly as the interference from the radar subsystem is jointly minimised. Comparing these K-IOP results with the non-optimised ZF baseline (cf. Fig.1) or the separate (radar-only or communication-only) waveform approaches in Sec. IX-BIX-C, it is evident that the integrated design offers substantially improved performance simultaneously for radar detection and communication reliability. This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
Figure 5: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
@@
Fig. 5 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for K-ROP, K-COP, and K-IOP optimisation techniques. At low $P_{\mathrm{R}}/N_{o}$ values (00-$10$ dB), K-IOP immediately achieves lower MSE than both K-ROP and K-COP. As $P_{\mathrm{R}}/N_{o}$ increases beyond 10 dB, the techniques begin to show more distinct performance. K-ROP demonstrates faster MSE decay, particularly in the $15$-$20$ dB range, while K-COP exhibits a steadier improvement but converges slightly more slowly. The K-IOP continues to outperform both separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values. Eventually, all three techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels. Notably, this MSE performance strongly correlates with the detection capabilities in Figs. 3.a, 2.a, and 4, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$. Conversely, missed detections or false alarms degrade the estimation by introducing noise-based or zero-valued entries.
Figure 6: Average CPU time (seconds) versus antenna array size $N$ for both optimisation techniques.
Fig. 6 presents the average CPU time versus antenna array size $N$, comparing all optimisation techniques. The K-ROP consistently demonstrates the lowest computational cost, with CPU times increasing from $0.001$ s at $N=10$ to $7.98$s at $N=100$. Meanwhile, K-COP exhibits moderate requirements, ranging from $0.04$ s to $14.07$s across the same $N$. By contrast, the K-IOP incurs the highest computational load, growing from $2.44$ s at $N=10$ to $37.69$s at $N=100$. For instance, at $N=50$, K-ROP requires only $2.80$s, whereas K-COP needs $3.90$s, and the K-IOP demands $19.96$s. These results indicate that while the radar-only and communication-only algorithms remain relatively efficient as $N$ increases, the K-IOP entails modestly higher computational complexity, while offering much higher performance.

### v5

@@ IX-B Radar Waveform KLD-based Optimisation
Figure 3: ISAC performance at $P_{\mathrm{r}}=0.5$ using two radar waveform optimisations: K-ROP and D-ROP. a) Radar subsystem: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 3 illustrates the performance of both subsystems using both KLD-based radar waveform optimisation (K-ROP) and detection probability radar waveform optimisation (D-ROP) techniques. In Fig. 3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ and detection probability $P_{D}$ versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance. The results in this figure also show that a slight improvement can be achieved using D-ROP compared to K-ROP, at the expense of slightly worse communication performance due to increased interference. Moreover, by comparing the results here with the ZF-based design in Fig. 1, it can be observed that a significant gain can be achieved using K-ROP and D-ROP. For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig. 1, while D-ROP achieves $\{97.4,90.7,99.99\}\%$, demonstrating improved detection capabilities by both optimisation techniques with D-ROP showing a slight improvement, at the cost of higher complexity. For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. While for D-ROP, $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ approximately $\{12.73,8.7,14.4\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB, increasing to $\{51.2,42.9,55.41\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system. The continued increase in $\mathrm{KLD}_{r}$ even as $P_{D}$ approaches unity demonstrates a key advantage of the KLD. While probability metrics saturate at their bounds, making it difficult to distinguish between detection probabilities closer to $1$, the unbounded KLD clearly differentiates these performance levels. This sensitivity is crucial for high-reliability applications and provides better numerical properties for optimisation algorithms.
@@
Fig. 3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ and BER versus $P_{\mathrm{R}}/N_{o}$. For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, respectively, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. D-ROP demonstrates similar trends with slightly lower performance, achieving $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values of $\{49.62,18.07,167.48\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB and $\{496.2,180.84,1674.79\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB for both techniques suggests they have reached their optimisation feasibility points, significantly enhancing the communication performance as the radar interference is effectively reduced. The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, while D-ROP reaches this threshold at $14$ dB, both significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted in Fig. 1. These results highlight the intricate relationship between radar detection capabilities and communication performance, with K-ROP demonstrating slightly superior performance over D-ROP.
@@ IX-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimisation is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$. The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
@@
Figure 4: ISAC performance at $P_{\mathrm{r}}=0.5$ using two communication waveform optimisations: K-COP and B-COP. a) Radar subsystem: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 4 illustrates the performance of both subsystems using both KLD-based communication waveform optimisation (K-COP) and BER-based communication waveform optimisation (B-COP) techniques. In Fig. 4.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ and detection probability $P_{D}$ versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. B-COP shows slightly lower $\mathrm{KLD_{\mathrm{r}}}$ performance, with values $\{9.31,7.48,10.04\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB, rising to $\{21.93,19.42,22.86\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform. At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while B-COP achieves $\{68.5,58.2,72.3\}\%$, compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig. 1. Both techniques demonstrate improved detection capabilities over the non-optimised scenario but remain below the radar waveform optimisation performance achieved using K-ROP in Fig. 3.
@@
Fig. 4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ and BER versus $P_{\mathrm{R}}/N_{o}$. For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. Interestingly, B-COP demonstrates marginally better communication performance. The BER performance reflects this trend, with B-COP achieving slightly better performance than K-COP. In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig. 1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs. B-COP achieves this threshold marginally earlier at $13.5$ dB. Both optimisation techniques provide significant performance gains over the non-optimised ZF, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec. IX-B due to radar interference, where K-COP’s fixed omnidirectional radar CIC matrix design creates an interference ceiling that limits achievable communication performance regardless of precoding sophistication. The communication performance shows less improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
@@ IX-D Integrated Waveform KLD-based Optimisation for ISAC
Figure 5: ISAC performance using the integrated waveform optimisation. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 5 illustrates the ISAC system performance under the KLD-based integrated waveform optimisation (K-IOP). In Fig. 5.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$. As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP. For instance, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, the achievable $\mathrm{KLD}_{\mathrm{r}}$ values for $\{\mathrm{T}_{1},\mathrm{T}_{2},\mathrm{T}_{3}\}$ are approximately $\{38.87,32.77,41.85\}$ bits, with corresponding detection probabilities $\{0.94,0.89,0.97\}$. Performance continues to rise at higher $\mathit{P}_{\mathrm{R}}/N_{o}$, achieving near-unity detection across all targets once $\mathit{P}_{\mathrm{R}}/N_{o}$ exceeds 22 dB.
@@
In Fig. 5.b, the communication subsystem performance is depicted via the BER and $\mathrm{KLD}_{\mathrm{c}}$ for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$. Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER. For example, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, $\mathrm{KLD}_{\mathrm{c}}$ values of $\{77.19,28.13,260.53\}$ bits are observed for $\{\mathrm{UE}_{1},\mathrm{UE}_{2},\mathrm{UE}_{3}\}$, with corresponding BERs of approximately $\{5.1\times 10^{-9},1.7\times 10^{-4},\lesssim 10^{-10}\}$. Beyond 14-16 dB, the BER for all UEs drops rapidly– often reaching effectively error-free communication– and $\mathrm{KLD}_{\mathrm{c}}$ grows significantly as the interference from the radar subsystem is jointly minimised. Comparing these K-IOP results with the non-optimised ZF baseline (cf. Fig.1) or the separate (radar-only or communication-only) waveform approaches in Sec. IX-BIX-C, it is evident that the integrated design offers substantially improved performance simultaneously for radar detection and communication reliability. This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
Figure 6: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$.
@@
Fig. 6 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques. At low $P_{\mathrm{R}}/N_{o}$ values (0-$10$ dB), K-IOP immediately achieves lower MSE than all of the other techniques. As $P_{\mathrm{R}}/N_{o}$ increases beyond $10$ dB, the techniques begin to show more distinct performance. K-ROP and D-ROP demonstrate faster MSE decay, particularly in the 1515-2020 dB range, while K-COP and B-COP exhibit a steadier improvement but converge slightly slower. The K-IOP continues to outperform all separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values. Eventually, all techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels. This MSE performance strongly correlates with the detection capabilities in Figs. 4.a, 3.a, and 5, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$. Conversely, missed detections or false alarms degrade the estimation by introducing noise-based estimates.
[…]
Fig. 7.a presents the average CPU time versus antenna array size $N$ for all optimisation techniques. The K-COP demonstrates the lowest computational cost, increasing from $0.04$s at $N=10$ to $7.98$s at $N=100$. On the other hand, K-ROP exhibits moderate requirements, ranging from $0.001$s to $14.07$s across the same $N$ range, and D-ROP shows higher CPU time growth, reaching $21.63$s at $N=100$. The communication-oriented B-COP requires $32.35$s at $N=100$, while K-IOP incurs the highest computational load at $37.69$s. These results confirm that while the radar-only optimisations maintain relatively low complexity, the joint K-IOP technique and B-COP require additional processing time, with K-IOP achieving superior ISAC performance despite its higher computational cost. Fig. 7.b illustrates the convergence behaviour of the $\mathrm{KLD_{ISAC}}$ metric versus the iteration count at $P_{\mathrm{R}}/N_{o}=18$ dB. As can be observed from the figure, K-IOP demonstrates the most rapid convergence, achieving substantial gains within the first $100$ iterations and reaching $74\%$ of its final value by iteration $158$. K-ROP also shows efficient convergence, stabilising at $95\%$ of its final value by iteration $135$, and D-ROP achieves stabilisation at $95\%$ of its final value by iteration $320$. The communication-focused technique K-COP converges quickly but saturates early, with K-COP reaching $97\%$ of its final value by iteration $90$, while B-COP exhibits a more gradual convergence pattern, requiring approximately $400$ iterations to reach its steady-state value. This analysis reveals that K-IOP efficiently balances rapid convergence with joint radar-communication optimisation, making it particularly suitable for real-time ISAC applications.
Figure 8: Achievable KLD under estimation errors for KLD-based optimisations, with $\sigma_{e}^{2}\!\in\!\{10^{-6},10^{-3},10^{-1}\}$. a) $\mathrm{KLD_{r}}$ $P_{\mathrm{R}}/N_{o}$. b) $\mathrm{KLD_{c}}$ vs $P_{\mathrm{R}}/N_{o}$.
Fig. 8 illustrates the effect of imperfect channel and radar target-response estimates on the KLD obtained by the proposed optimisation techniques. For clarity, this is shown for $\mathrm{UE}_{1}$, and $\mathrm{T}_{1}$. Fig. 8.a, presents $\mathrm{KLD_{r}}$ versus $P_{\mathrm{R}}/N_{o}$, and Fig. 8.b presents $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for three error-variance levels. For each optimisation technique, the quasi-perfect case ($\sigma_{e}^{2}=10^{-6}$) yields the highest KLD; increasing the error variance to $10^{-3}$ noticeably flattens the curves, and at $10^{-1}$ the gains from all schemes almost disappear. As the error variance rises from $10^{-6}$ to $10^{-1}$, both the radar and communication KLD curves flatten, producing a clear error floor at high $P_{\mathrm{R}}/N_{o}$. The observed degradation is therefore a predictable consequence of reduced CSI fidelity, underlining the importance of accurate estimation.